40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Feb 17 23:50:55 2011, tara, DailyProgress, VCO, PDH box irregular gain behavior PDH_input_refernoise.pngACAVTF.pngnoise.png
    Reply  Fri Feb 18 23:52:42 2011, tara, DailyProgress, VCO, PDH box irregular gain behavior VCOTF.pngrefernoise.pngbeat_2011_02_18.png
Message ID: 504     Entry time: Fri Feb 18 23:52:42 2011     In reply to: 503
Author: tara 
Type: DailyProgress 
Category: VCO 
Subject: PDH box irregular gain behavior 

The input referred noise for the PDH box is measured and corrected.  The input referred noise (IRN) of the PDH box

is getting lower from gain 4 to 5 and increase at gain 6 and does not change until gain 10(max).

The IRN (from PDH + RFPD setup)is converted to laser frequency noise and put on the noise budget.

It seems to be a limiting source at high frequency.



Frank pointed out that I made a mistake my measuring the whole OLGTF of ACAV instead of the PDH' TF only.

So I remeasure the PDH's TF at different gain. For the range of interest, the TF has a pole at 48 Hz. The TF's DC gain for each gain settings are

gain   DC gain [dB]
4        55
5        60
6        65
7        70
8        75
9        80
10       85

The TF for each gains and the fit are plot below, they are offset for comparison. (The name should be PDH TF file.)


The input referred noise is then calculated, and plotted below (fig2)

For PDH alone (Line plots), the input referred noise (PDH alone) at gain 5 is smallest. gain 4 is a bit higher, gain 6 to 10 are about the same. 

For PDH + RFPD setup (dot lines), the input referred noise at gain 4 and 5 are about the same,

and getting higher at gain 6  and remains  the same to gain 10.


The IRN (PDH + RFPD set) from gain 5 and 7 are chosen for the noise budget plot as our gain is always set around this.

(I will interpolate it to lower frequency with 1/f slope, the actual data will be measured)


 To convert it to frequency noise, the IRN is multiplied by the slope of the error signal, 4.67 MHz/V. (See below for detail.)

and plot it together on the noise budget.


This time, the beat signal (green)is measured with smaller input range (10kHz),

so the signal is not limited by LO phase noise and goes below the noise budget .

However, the IRN (from PDH+ RFPD) noise at gain 7 is higher than the beat noise, by a factor of 2, but the IRN at gain 5 

matches the signal. The peaks from IRN and beat b/w 100 to 1kHz matches quite well

Note that LO phase noise from 10kHz input range is l~ 30-40 mHz at this BW which is lower than the measured beat noise.

(I'll find the data for LO phase noise and plot it together





error signal slope msmt


I scanned the laser by injecting a sinusoidal signal via fast channel. (Unlock laser from RCAV.)

Set  SLOWDC so the frequency is close to ACAV resonance, and measure the signal from the mixer output.

The pk - pk is 22.8 mV. The frequency between pk-pk is FWHM of the cavity = 2x ACAV's pole = 108 kHz,

so the slope is 108 kHz/ 22.8 mV = 4.74 [MHz/V]



I measured the input referred noise of the VCO's PDH box at different gain setting.

The input refer noise varies with gain setup. The result is plotted below


When both cavities are in lock, the VCO receives the feedback signal from the PDH box to control the laser frequency.

The PSD of the feedback signal to VCO varies with gain setup on the PDH box significantly, so it needs to be checked.

To decide the gain setup range for test, the gain is adjusted while the beat noise is observed. The minimum gain and maximum gain

where the beat noise still looks nice are chosen, which are 4 and 7 as read on the knob dial.


Then the noise of the PDH box is measured at different gain setups (4,5,6,7.) When

1) the PDH input is shorted (for PDH intrinsic noise alone ), and

2) the beam on the RFPD is blocked (for whole set of RFPD, mixer and PDH box)

 (data from gain 4 with RFPD has to be remeasured )


Then the open loop TF is measured. The measured noise divided by the TF is the input referred noise of our system.

The TF in the measured region can be approximated with 1/f slope.

 The input referred goes down alot from gain 4 to 5, and not much from 5 to 7;


The input referred noise doesn't seem correct. I'll double check it later.




Attachment 1: VCOTF.png  20 kB  | Hide | Hide all
Attachment 2: refernoise.png  64 kB  | Hide | Hide all
Attachment 3: beat_2011_02_18.png  51 kB  | Hide | Hide all
ELOG V3.1.3-